F{x}=2x squared
when x equals -3
find f{-3}

Answer by sdmmadam@yahoo.com(530)

F{x}=2x squared
when x equals -3
find f{-3}
It should be find f(-3)
F(x) = 2x squared
Does it mean (2x) squared or 2times(x squared)?
Please ascertain from your teacher what exactly 2x squared means.
We hail from two different nations you see!.The interpretations might vary.
Let me do both the versions for you.
1 ) F(x) = (2x)^2
That is F(x) = 2^2x^2 = 4(x^2)
Putting x = -3
F(-3) = 4X[(-3)^2]
= 4X9 = 36
2) F(x) = 2(x^2)
F(-3) = 2X[(-3)^2]
= 2X9 = 18